OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x+y=7 9x+[ ]y=[ ]

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

If they are coincident, than it means that they are actually the same equation

OpenStudy (anonymous):

\[ \to 3x+y=7\\ \to 9x+[ ]y=[ ] \]It looks like the second is just 3 times the first.

OpenStudy (anonymous):

So \(3\cdot y=3y\) and \(3\cdot 7=21\)

OpenStudy (anonymous):

ok so what about the middle one?